# One-Step Offline Distillation of Diffusion-based Models via Koopman Modeling

### v2

@@ E.1 Training and Sampling Procedures Psudo-codes
1:Training pairs $(x_{0},x_{T})$, labels $y$ (optional), model $E_{\phi}$, $E_{\psi}$,$C_{\eta}$, $D_{\psi}$ discriminator $D_{\gamma}$
@@
13:     Compute reconstruction loss: $\mathcal{L}_{\text{rec}}\leftarrow d(x_{0},\hat{x_{0}})$
@@ E.2 Datasets Collection
For the Checkerboard dataset, we sampled using 10 NFEs with EDM [28] or flow matching [40]. Following [19], we sampled CIFAR-10 with 35 NFEs from EDM. For FFHQ and AFHQv2, we used 79 NFEs using EDM, as recommended.
@@ E.3 Method Implementation Details
For this dataset, we build upon the implementation of [41]. Both $x_{0}$ and $x_{T}$ are encoded using the same denoising architecture as described in their work. The decoding of $z_{0}$ is performed using the inverse of this module. The Koopman operator is implemented as a simple linear layer. Complete implementation details are available in the shared code.
@@
We adopt the SongUNet architecture from [28] for encoding $x_{0}$ and $x_{T}$, as well as for decoding $z_{0}$. To reduce latent dimensionality, the output channels of the network are set to 1, effectively reducing the latent space size by a factor of 3. The Koopman operator is implemented either as a standard linear transformation or using our custom Fast Koopman module. To maintain architectural fairness, we downscale all networks to match the total parameter count of the original Karras network [28] and the GET distillation model [19]. The FFHQ and AFHQv2 configurations follow a similar encoder-decoder structure with one key difference: a linear projection layer is appended to the end of the encoder and the beginning of the decoder to further control the latent space dimensionality. All other aspects remain consistent across datasets. Note that the discriminator is trained using a separate optimizer, akin to the Koopman module’s optimizer.
@@ E.4 Human Evaluation Experiment
While the Fréchet Inception Distance (FID) is widely used to evaluate the quality of images generated by generative models, recent studies have highlighted its limitations in aligning with human judgments of image realism. For instance, Jayasumana et al. [27] demonstrate that FID can contradict human evaluations, particularly in assessing perceptual quality and semantic alignment, due to its reliance on Inception-v3 features and the assumption of Gaussian distributions. Similarly, Otani et al. [54] emphasize that automatic metrics like FID often fail to capture the nuanced aspects of human perception, underscoring the need for standardized human evaluation protocols. Furthermore, Kynkäänniemi et al. [38] reveal that FID can be manipulated by aligning top-N class histograms without genuinely improving image quality, indicating potential biases in the metric. These findings suggest that while FID provides a useful quantitative signal, it may not fully encapsulate human perceptions of image realism, highlighting the importance of incorporating human evaluations when assessing generative models.
@@
Therefore, in the main paper, we report the results of a human evaluation experiment designed to compare our method against EDM. Specifically, we curated 10 image collages (6$\times$6) of human faces generated by the teacher model (EDM, 79 NFEs) and 10 collages generated by our student model (1 NFE), using samples from the FFHQ dataset. Each comparison slide presents one collage from each model, displayed side-by-side as illustrated in Fig. 4, with the order randomized. We then asked 30 human raters to evaluate 10 such slides and choose, for each, which collage appeared more realistic. If both appeared equally realistic, they were instructed to make a random choice. This process yielded 300 total decisions. We report in the main paper and in Tab. 8, the proportion of times each model was preferred. Finally, we attach to the appendix both the slides that we sent to the practitioners and the Excel that aggregates the results. The results indicate that decisions were made almost at random, suggesting that participants had difficulty distinguishing between the student and teacher generations. This implies that, despite existing FID gaps, human evaluators found it hard to perceive meaningful differences. Participants reported that both collages included highly realistic images, as well as images with noticeable artifacts that appeared unrealistic, in their judgment.
E.5 Parameter Scaling
@@
E.6 Extended Outlier Analysis
@@ Appendix F Additional Discussion of Future Work and Current Limitations
The connection between Koopman theory and diffusion model distillation opens the door to a rich body of theoretical and practical tools. On the theoretical side, works on sample complexity and error bounds for Koopman operators [56, 10, 51] may provide insights into convergence guarantees. Information-theoretic perspectives [42] could offer new modeling constraints or regularization schemes. On the practical side, methods for rare-event sampling [76] may improve long-tail generation capabilities. Finally, while the current mathematical formulation supports discrete Koopman evaluation, the underlying theory naturally extends to continuous settings as well. Exploring these directions could substantially extend and generalize the impact of our framework.
@@
While adversarial components have been shown to be critical in many distillation methods [74, 32], our current implementation uses a basic discriminator architecture with standard loss functions and training schemes. We believe that integrating more advanced adversarial paradigms could significantly boost performance, potentially closing the gap with state-of-the-art online distillation methods.
@@
Due to limited computational resources, we were unable to extend our offline distillation framework to more complex tasks, such as text-to-image generation, that require large-scale compute infrastructures. It will be interesting to explore how our method can scale or adapt to other domains and tasks. Additionally, because our method makes no assumptions about the initial and terminal states of the underlying dynamical system, we believe it is applicable to a broader class of generative models, including DDBMs [80], stochastic interpolants [1], and others [13].

### v3

@@ E.1 Training and Sampling Procedures Psudo-codes
1:Training pairs $(x_{0},x_{T})$, labels $y$ (optional), model $E_{\phi}$, $E_{\psi}$$C_{\eta}$, $D_{\psi}$ discriminator $D_{\gamma}$
@@
[…]
@@ E.2 Datasets Collection
For the Checkerboard dataset, we sampled using 10 NFEs with EDM karras2022elucidating or flow matching lipman2022flow . Following geng2023one , we sampled CIFAR-10 with 35 NFEs from EDM. For FFHQ and AFHQv2, we used 79 NFEs using EDM, as recommended.
@@ E.3 Method Implementation Details
For this dataset, we build upon the implementation of lipman2023flow . Both $x_{0}$ and $x_{T}$ are encoded using the same denoising architecture as described in their work. The decoding of $z_{0}$ is performed using the inverse of this module. The Koopman operator is implemented as a simple linear layer. Complete implementation details are available in the shared code.
@@
We adopt the SongUNet architecture from karras2022elucidating for encoding $x_{0}$ and $x_{T}$, as well as for decoding $z_{0}$. To reduce latent dimensionality, the output channels of the network are set to 1, effectively reducing the latent space size by a factor of 3. The Koopman operator is implemented either as a standard linear transformation or using our custom Fast Koopman module. To maintain architectural fairness, we downscale all networks to match the total parameter count of the original Karras network karras2022elucidating and the GET distillation model geng2023one . The FFHQ and AFHQv2 configurations follow a similar encoder-decoder structure with one key difference: a linear projection layer is appended to the end of the encoder and the beginning of the decoder to further control the latent space dimensionality. All other aspects remain consistent across datasets. Note that the discriminator is trained using a separate optimizer, akin to the Koopman module’s optimizer.
@@ E.4 Human Evaluation Experiment
While the Fréchet Inception Distance (FID) is widely used to evaluate the quality of images generated by generative models, recent studies have highlighted its limitations in aligning with human judgments of image realism. For instance, Jayasumana et al. jayasumana2024rethinking demonstrate that FID can contradict human evaluations, particularly in assessing perceptual quality and semantic alignment, due to its reliance on Inception-v3 features and the assumption of Gaussian distributions. Similarly, Otani et al. otani2023toward emphasize that automatic metrics like FID often fail to capture the nuanced aspects of human perception, underscoring the need for standardized human evaluation protocols. Furthermore, Kynkäänniemi et al. kynkaanniemi2022role reveal that FID can be manipulated by aligning top-N class histograms without genuinely improving image quality, indicating potential biases in the metric. These findings suggest that while FID provides a useful quantitative signal, it may not fully encapsulate human perceptions of image realism, highlighting the importance of incorporating human evaluations when assessing generative models.
@@
Therefore, in the main paper, we report the results of a human evaluation experiment designed to compare our method against EDM. Specifically, we curated 10 image collages ($\times6$) of human faces generated by the teacher model (EDM, 79 NFEs) and 10 collages generated by our student model (1 NFE), using samples from the FFHQ dataset. Each comparison slide presents one collage from each model, displayed side-by-side as illustrated in Fig. 4, with the order randomized. We then asked 30 human raters to evaluate 10 such slides and choose, for each, which collage appeared more realistic. If both appeared equally realistic, they were instructed to make a random choice. This process yielded 300 total decisions. We report in the main paper and in Tab. 8, the proportion of times each model was preferred. Finally, we attach to the appendix both the slides that we sent to the practitioners and the Excel that aggregates the results. The results indicate that decisions were made almost at random, suggesting that participants had difficulty distinguishing between the student and teacher generations. This implies that, despite existing FID gaps, human evaluators found it hard to perceive meaningful differences. Participants reported that both collages included highly realistic images, as well as images with noticeable artifacts that appeared unrealistic, in their judgment.
E.5 Cost and Efficiency of Offline Distillation
An important consideration in offline distillation is the total computational cost, which includes both the training of the student model and the one-time cost of generating the teacher-produced noise–data pairs. While this pair generation step incurs an initial overhead, it is performed only once and can be reused across multiple student architectures or experiments, effectively amortizing its cost. Following the analysis of (geng2023one, ), which explicitly incorporates the cost of data-pair generation into its complexity assessment, the overall computational load of offline distillation remains favorable compared to online approaches. This is primarily because online distillation requires repeated access to the full teacher model throughout training, resulting in substantially higher neural function evaluations and memory usage. In our experiments on the main paper, we observe that even when accounting for the data generation step, our method achieves a markedly better trade-off between cost and performance than prior offline frameworks such as GET. These results collectively support that, despite the initial preprocessing overhead, offline distillation constitutes an efficient and scalable alternative for practical deployment. To see the full analysis between online and offline approaches, please refer to geng2023one .
E.6 Parameter Scaling
@@
E.7 Extended Outlier Analysis
@@ Appendix F Additional Discussion of Future Work and Current Limitations
The connection between Koopman theory and diffusion model distillation opens the door to a rich body of theoretical and practical tools. On the theoretical side, works on sample complexity and error bounds for Koopman operators philipp2023error ; chen2018sample ; muthirayan2021online may provide insights into convergence guarantees. Information-theoretic perspectives lozano2022information could offer new modeling constraints or regularization schemes. On the practical side, methods for rare-event sampling zhang2022koopman may improve long-tail generation capabilities. Finally, while the current mathematical formulation supports discrete Koopman evaluation, the underlying theory naturally extends to continuous settings as well. Exploring these directions could substantially extend and generalize the impact of our framework.
@@
While adversarial components have been shown to be critical in many distillation methods yin2024improved ; kim2024consistency , our current implementation uses a basic discriminator architecture with standard loss functions and training schemes. We believe that integrating more advanced adversarial paradigms could significantly boost performance, potentially closing the gap with state-of-the-art online distillation methods.
@@
Due to limited computational resources, we were unable to extend our offline distillation framework to more complex tasks, such as text-to-image generation, that require large-scale compute infrastructures. It will be interesting to explore how our method can scale or adapt to other domains and tasks. Additionally, because our method makes no assumptions about the initial and terminal states of the underlying dynamical system, we believe it is applicable to a broader class of generative models, including DDBMs zhou2023denoising , stochastic interpolants albergo2023stochastic , and others de2021diffusion .